Ten students apply for the position of grader in mathematics. one grader is to be assigned to each of four teachers. In how many different ways can the assignment of graders be if no student grades for two teachers?
Answer by Aldorozos(172)

It is a permutation. Let's assum we have four seats and we have 12 students. How many ways these students can be seated.
One way is 1234, another is 1243, another 1235 and so on.
The formula for this is n!/(n-r)!
12!/(12-4)! = 12*11*10*9*8!/8!= 12*11*10*9